OpenStudy (anonymous):

graph arg(Z+i)-arg(Z-i)=pi/4

OpenStudy (amriju):

u can remember hundreds of special cases...but u have a general...but a bit lenghty rule....take z=x+iy....now put it in the eqn....u know that arg( a+ib)=tan^-1(b/a)...use this defination....then solve for tan^-1((y+1)/x)-tan^-1((y-1)/x)=tan^-1(1)...u get the relation betn x and y...thats ur locus...essentially it will be a circle...centre probb...(1,0) with radius 1.4

OpenStudy (anonymous):

hey im not really sure how to simplify after the tan^-1 part..

OpenStudy (anonymous):

@amriju help!!

OpenStudy (amriju):

u kno the formula...? of \[\tan^{-1} a-\tan^{-1} b\]

OpenStudy (anonymous):

noep

OpenStudy (anonymous):

how to combine

OpenStudy (amriju):

\[\tan^{-1} a-\tan^{-1} b=\tan^{-1} ((a-b)/(1+ab))\]..but u hav to be good in trigo to solve complex numbers...u kno..?

OpenStudy (anonymous):

nope.. dont think i've learnt this before

OpenStudy (amriju):

u HAVE TO kno dis...else u wont be able to work out problems on arguments

OpenStudy (anonymous):

oh.. so cant do other wya?

OpenStudy (amriju):

yeah....there r different cases lyk..." if this is given, the locus will be this..."...learn all of them...

OpenStudy (anonymous):

yeah i cant do these.. trouble with arg ones

OpenStudy (amriju):

thats coz ur trigo is weak...

OpenStudy (anonymous):

how so?? i've never learnt much about the inverse trig functions

OpenStudy (amriju):

u betr learn...to solve arg u need it

OpenStudy (anonymous):

how do i lear??

OpenStudy (anonymous):

can you teach me

OpenStudy (amriju):

its a pretty large chapter...u shud look for a maths specialist....